# Beyond Sparsity: Quantum Block Encoding for Dense Matrices via Hierarchically Low Rank Compression

### v2

@@ Abstract
While quantum algorithms for solving large scale systems of linear equations offer potential speedups, their application has largely been confined to sparse matrices. This work extends the scope of these algorithms to a broad class of structured dense matrices arise in potential theory, covariance modeling, and computational physics, namely, hierarchically block separable (HBS) matrices. We develop two distinct methods to make these systems amenable to quantum solvers. The first is a pre-processing approach that transforms the dense matrix into a larger but sparse format. The second is a direct block encoding scheme that recursively constructs the necessary oracles from the HBS structure. We provide a detailed complexity analysis and rigorous error bounds for both methods. Numerical experiments are presented to validate the effectiveness of our approaches.
I Introduction
@@
However, the performance of these algorithms relies on a key assumption: sparsity. The HHL algorithm, for example, depends on the ability to efficiently simulate the Hamiltonian evolution of the system matrix. The efficiency of this simulation, in turn, depends on the matrix’s sparsity [childsFirstQuantumSimulation2018]. Similarly, QSVT-based approaches require access to a block encoding of the system matrix. While constructing a block encoding for a sparse matrix is relatively straightforward, doing so for a dense matrix remains a challenge. This sparsity issue becomes severe because, under data access models similar to those used for block encodings, it is possible to construct classical randomized algorithms that match the performance of their quantum counterparts [gilyenImprovedQuantuminspiredAlgorithm2022]. As a result, better approaches for dense matrices are needed, but existing methods have their own drawbacks. For instance, the quantum algorithms studied in [wossnigQuantumLinearSystem2018] handle dense matrices with only a polynomial speedup instead of an exponential one.
@@
Motivated by the work of Ho and Greengard [hoFastDirectSolver2012], we proposes and analyzes two quantum approaches for solving dense linear systems based on hierarchical matrix representations. The first is an extended sparsification method that transforms the dense matrix into a larger, sparse system, making it amenable to efficient block encoding designed for sparse matrices. The second is a direct, recursive construction of the block encoding that follows the hierarchical structure. We provide a rigorous analysis for both strategies. For the sparsification approach, we show that the construction cost and resulting sparsity scale linearly with the matrix dimension. The condition number and approximation errors are also provided. For the recursive approach, we derive explicit bounds on the growth of the subnormalization factor and the accumulation of precision error.
@@
The remainder of this paper is organized as follows. In Section 2, we review the basic formulations of quantum linear system algorithms and show the obstacles in constructing block encodings for dense matrices. Section 3 introduces HBS matrices and presents a modified algorithm adapted for the subsequential block encoding in quantum setting. In Section 4, we propose a sparsification based quantum approach, which uses an extended sparsification to combine the HBS matrix components into a single but slightly larger system. As an alternative, Section 5 describes a direct method for constructing the block encoding from the hierarchical matrix structure. Section 6 presents numerical results from experiments on integral equations to evaluate the performance of our methods based on the classical computers. Finally, Section 7 provides concluding remarks and discusses potential directions for future work.
II quantum linear system algorithms
@@
Definition II.2 (Block Encoding).
@@
Lemma II.1 (Block encoding [nguyenBlockencodingDenseFullrank2022]).
@@
Lemma II.1 establishes that the efficiency of the block encoding, specifically the subnormalization factor $\alpha$, is determined by the geometric mean of the row and column sparsities. While this construction is highly efficient for sparse matrices, its limitations become apparent for dense matrices. For example, if $s_{r}=s_{c}=2^{n}$, the subnormalization factor $\alpha=\hat{a}\sqrt{s_{r}s_{c}}$ scales as $\hat{a}\cdot 2^{n}$. Since the complexity of algorithms using this block encoding typically scales polynomially with $\alpha$, this exponential growth in $n$ poses a significant drawback. We address this problem by proposing and analyzing two methods. The first uses an extended sparsification technique to convert the dense matrix into a sparse one, followed by an analysis of the cost to construct a block encoding for the resulting system. The second is a direct method for block encoding the hierarchical matrix structure itself, for which we analyze the complexity and the required subnormalization factor.
In the complexity analysis that follow, we account for the number of non-zero elements. This number provides a more realistic measure of resource requirements than simply using the matrix dimensions. The cost of implementing the oracles, such as $\mathcal{O}_{r},\mathcal{O}_{c}$ and $\mathcal{O}_{A}$ in Lemma II.1, is directly related to the gate count needed to access the value and location of each non-zero entry. In particular, [zhangCircuitComplexityQuantum2024] shows that the gate count for constructing these oracles is directly proportional to the number of non-zero elements in the matrix representation. Therefore, to implement our proposed methods effectively, we first require a suitable way to compress the dense matrix into a format with a manageable number of entries. The next section introduces the hierarchically block separable (HBS) format [hoFastSemidirectLeast2014] that serves as the foundation for our analysis.
III Hierarchically block separable matrices
Solving large scale linear systems is a classical problem in scientific computing. While standard iterative methods like Bi-CGSTAB and GMRES are often applied, their computational efficiency is matrix dependent [doi:10.1137/1.9781421407944], which can lead to poor performance for the dense matrices that arise in many physical problems. An effective strategy is to exploit the specific structure hidden behind these matrices. For instance, matrices from the discretization of integral equations are not arbitrarily dense but possess a sparse structure. This property stems from the smoothness of the underlying kernel $K(x,y)$ for well separated points $x$ and $y$ in 2D or 3D. Important kernels in potential theory and wave physics that exhibit this behavior include:
@@ III Hierarchically block separable matrices
This recursive compression algorithm, illustrated in Figure 1, constructs the hierarchical factors ($D_{l},L_{l},R_{l}$) level by level. Let $S_{0}=A$ be the initial matrix. At each level $l$ of the hierarchy, the algorithm processes the current matrix $S_{l-1}$, which is partitioned into blocks based on the spatial tree. The blocks corresponding to near field, where the target $x$ and source $y$ are close, are typically dense, and these blocks are extracted and collectively form the diagonal factor $D_{l}$. The remaining matrix, which we denote as $B_{l}=S_{l-1}-D_{l}$, contains only the far field interactions. These far field blocks are presumed to be numerically low-rank. They are approximated using ID to identify a small set of skeleton indices and to construct basis matrices. To perform this efficiently, all far-field blocks within a given row partition share a common basis matrix, and similarly, all blocks within a column partition share a common column basis. The computational cost of these ID procedures can be significantly reduced by using proxy approximation, which replaces global computations with local ones involving a small set of sources [xingInterpolativeDecompositionProxy2020]; a more detailed description is provided in Appendix LABEL:ap:classic. These bases are then assembled into the block-diagonal factors $L_{l}$ and $R_{l}$. The far field part is thus approximated as $B_{l}\approx L_{l}S_{l}R_{l}$, where $S_{l}$ is the skeleton matrix. This new matrix $S_{l}$ is passed to the next level, $l+1$, for the same decomposition procedure. This process is repeated until a predefined recursion depth $\lambda$ is reached, and the final remaining skeleton matrix is set as $D_{\lambda+1}$.
@@
To address this limitation, based on the HBS algorithm [hoFastDirectSolver2012], a more robust strategy uses the geometric coordinates of the underlying data points, which are typically available for matrices defined by a kernel function [coulierInverseFastMultipole2017]. Instead of partitioning indices based on their ordering, the hierarchy is constructed directly from these spatial coordinates, for instance, by employing a quadtree in 2D or an octree in 3D. For matrices that are not strictly generated by a kernel function, precise coordinates are not strictly required for this partitioning; approximate coordinates are sufficient. For example, if the matrix represents a kernel-based system subject to a small perturbation, the coordinates associated with the original kernel function can still be effectively used. This hierarchical structure ensures that matrix blocks corresponding to well-separated spatial clusters are low rank and compressible, whereas blocks corresponding to adjacent points are generally treated as dense diagonal blocks at a given level. The procedure is summarized in Algorithm III and illustrated in Figure 1.
@@
non-zero elements.
The proof of Theorem III.1 can be found in Appendix LABEL:ap:classic. This theorem shows that the classical preprocessing cost and the number of non-zero elements in the resulting matrices both scale linearly with the matrix dimension $N$. Furthermore, the entries of these matrices are bounded. These properties are important considerations for quantum algorithms, and we will apply the results of Theorem III.1 in the analysis that follows.
IV A Sparsification Based Quantum Approach

### v3

@@ Abstract
While quantum algorithms for solving large scale systems of linear equations offer potentially exponential speedups, their application has largely been confined to sparse matrices. This work extends the scope of these algorithms to a broad class of structured dense matrices arise in potential theory, covariance modeling, and computational physics, namely, hierarchically block separable (HBS) matrices. We develop two distinct methods to make these systems amenable to quantum solvers. The first is a pre-processing approach that transforms the dense matrix into a larger but sparse format. The second is a direct block encoding scheme that recursively constructs the necessary oracles from the HBS structure. We provide a detailed complexity analysis and rigorous error bounds for both methods. Numerical experiments are presented to validate the effectiveness of our approaches.
1 Introduction
@@
However, the performance of these algorithms relies on a key assumption: sparsity. The HHL algorithm, for example, depends on the ability to efficiently simulate the Hamiltonian evolution of the system matrix. The efficiency of this simulation, in turn, depends on the matrix’s sparsity [childsFirstQuantumSimulation2018]. Similarly, QSVT-based approaches require access to a block encoding of the system matrix. While constructing a block encoding for a sparse matrix is relatively straightforward, doing so for a dense matrix remains a challenge. This sparsity issue becomes severe because, under data access models similar to those used for block encodings, it is possible to construct classical randomized algorithms that match the performance of their quantum counterparts [gilyenImprovedQuantuminspiredAlgorithm2022]. Consequently, better approaches for dense matrices are needed, but existing methods have their own drawbacks. For instance, the quantum algorithms studied in [wossnigQuantumLinearSystem2018] handle dense matrices with only a polynomial speedup instead of an exponential one.
@@
1.1 Our contributions
We formally analyze the compressibility of dense matrices under the hierarchically block separable (HBS) framework [hoFastDirectSolver2012]. We quantify how such matrices can be reduced to linear size with uniformly bounded entries. In particular, proxy compression and strong RRQR-based interpolative decomposition enable the construction of hierarchical factors with $O(N)$ computational cost and $O(N)$ memory size. The precise statement are given in Theorem 3.1.
Based on this hierarchical structure, we introduce a quantum algorithm that uses sparsification. We combine the hierarchical factors into a single extended sparse system whose dimension grows only linearly, while its row and column sparsities remain bounded. This procedure makes the system compatible with standard sparse matrix block encoding oracles.
Main Result 1.
Let $A_{\mathrm{sp}}$ be the extended sparse system derived from an HBS approximation of a dense matrix $A$. Then $A_{\mathrm{sp}}$ has bounded row and column sparsities and contains only $O(N)$ nonzero entries. Consequently, one can implement a block encoding of $A_{\mathrm{sp}}$ with subnormalization factor bounded by a constant (independent of $N$), using oracles that access only $O(N)$ stored nonzeros.
A detailed formulation in the oracle model is given in Theorem 4.1.
We also analyze the side effects introduced by extended sparsification. Our analysis addresses the probability loss from auxiliary variables, the conditional number of the system via a regularized formulation and the propagation of the HBS approximation error to the solution. The precise statement and constants are given in Theorem 4.2.
Finally, we develop a direct recursive construction of a block encoding for the HBS approximation itself. By using QMM to multiply block encodings along the hierarchy and LCU to add near field blocks level by level, this construction yields upper bounds for subnormalization growth and error accumulation.
Main Result 2.
Given block encodings of the HBS factors $\{D_{\ell},L_{\ell},R_{\ell}\}$ with bounded subnormalization factors, there is a recursive procedure (based on QMM and LCU) that produces a block encoding of the full HBS approximation. The resulting subnormalization factor and overall precision loss admit upper bounds determined by the hierarchy depth and the factor-wise subnormalizations.
The complete quantitative statement is Theorem LABEL:thm:recursive_be_analysis in Section 5. A detailed derivation of the subnormalization and error bounds is provided immediately after the theorem.
1.2 Related works and comparison
The concept of using hierarchical matrix structures for quantum block encoding was recently explored by Nguyen et al. [nguyenBlockencodingDenseFullrank2022]. They introduced a method for block encoding dense kernel matrices by splitting them into block sparse hierarchical levels and using a linear combination of unitaries (LCU). Their approach achieves optimal subnormalization factors for kernels with specific analytical decay properties. However, because it evaluates entries across all dense admissible blocks at each level, the strategy’s data access cost scales as $O(N^{2})$ for general matrices.
Our work takes a different approach by using the HBS format. We employ nested interpolative decompositions and proxy surfaces to compress the matrix into sparse factors with only $O(N)$ nonzeros. Furthermore, our method removes the need for an analytical kernel representation, making it applicable to a broader range of structured or perturbed matrices.
@@ 1.2 Related works and comparison
The remainder of this paper is organized as follows. In Section 2, we review the basic formulations of quantum linear system algorithms and show the obstacles in constructing block encodings for dense matrices. Section 3 introduces HBS matrices and presents a modified algorithm adapted for the subsequential block encoding in quantum setting. In Section 4, we propose a sparsification based quantum approach, which uses an extended sparsification to combine the HBS matrix components into a single but slightly larger system. As an alternative, Section 5 describes a direct method for constructing the block encoding from the hierarchical matrix structure. Section LABEL:se:numerical presents numerical results from experiments on integral equations to evaluate the performance of our methods based on the classical computers. Finally, Section LABEL:se:conclusions provides concluding remarks and discusses potential directions for future work.
2 Quantum linear system algorithms
@@
Definition 2.2 (Block Encoding).
@@
Lemma 2.1 (Block encoding [nguyenBlockencodingDenseFullrank2022]).
@@
Lemma 2.1 establishes that the efficiency of the block encoding, specifically the subnormalization factor $\alpha$, is determined by the geometric mean of the row and column sparsities. While this construction is highly efficient for sparse matrices, its limitations become apparent for dense matrices. For example, if $s_{r}=s_{c}=2^{n}$, the subnormalization factor $\alpha=\hat{a}\sqrt{s_{r}s_{c}}$ scales as $\hat{a}\cdot 2^{n}$. Since the complexity of algorithms using this block encoding typically scales polynomially with $\alpha$, this exponential growth in $n$ poses a significant drawback. We address this problem by proposing and analyzing two methods. The first uses an extended sparsification technique to convert the dense matrix into a sparse one, followed by an analysis of the cost to construct a block encoding for the resulting system. The second is a direct method for block encoding the hierarchical matrix structure itself, for which we analyze the complexity and the required subnormalization factor.
In the complexity analysis that follow, we account for the number of nonzero elements. This number provides a more realistic measure of resource requirements than simply using the matrix dimensions. The cost of implementing the oracles, such as $\mathcal{O}_{r},\mathcal{O}_{c}$ and $\mathcal{O}_{A}$ in Lemma 2.1, is directly related to the gate count needed to access the value and location of each nonzero entry. In particular, [zhangCircuitComplexityQuantum2024] shows that the gate count for constructing these oracles is directly proportional to the number of nonzero elements in the matrix representation. Therefore, to implement our proposed methods effectively, we first require a suitable way to compress the dense matrix into a format with a manageable number of entries. The next section introduces the hierarchically block separable (HBS) format [hoFastSemidirectLeast2014] that serves as the foundation for our analysis.
3 Hierarchically block separable matrices
Solving large scale linear systems is a classical problem in scientific computing. While standard iterative methods like Bi-CGSTAB and GMRES are often applied, their computational efficiency is matrix dependent [doi:10.1137/1.9781421407944], which can lead to poor performance for the dense matrices that arise in many physical problems. An effective strategy is to exploit the specific structure hidden behind these matrices. In particular, matrices from the discretization of integral equations are not arbitrarily dense but possess a sparse structure. This property stems from the smoothness of the underlying kernel $K(x,y)$ for well separated points $x$ and $y$ in 2D or 3D. Important kernels in potential theory and wave physics that exhibit this behavior include:
@@ 3 Hierarchically block separable matrices
This recursive compression algorithm, illustrated in Figure 1, constructs the hierarchical factors ($D_{l},L_{l},R_{l}$) level by level. Let $S_{0}=A$ be the initial matrix. At each level $l$ of the hierarchy, the algorithm processes the current matrix $S_{l-1}$, which is partitioned into blocks based on the spatial tree. The blocks corresponding to near field, where the target $x$ and source $y$ are close, are typically dense, and these blocks are extracted and collectively form the diagonal factor $D_{l}$. The remaining matrix, which we denote as $B_{l}=S_{l-1}-D_{l}$, contains only the far field interactions. These far field blocks are presumed to be numerically low-rank. They are approximated using ID to identify a small set of skeleton indices and to construct basis matrices. To perform this efficiently, all far-field blocks within a given row partition share a common basis matrix, and similarly, all blocks within a column partition share a common column basis. The computational cost of these ID procedures can be significantly reduced by using proxy approximation, which replaces global computations with local ones involving a small set of sources [xingInterpolativeDecompositionProxy2020]; a more detailed description is provided in Section 3.1. These bases are then assembled into the block-diagonal factors $L_{l}$ and $R_{l}$. The far field part is thus approximated as $B_{l}\approx L_{l}S_{l}R_{l}$, where $S_{l}$ is the skeleton matrix. This new matrix $S_{l}$ is passed to the next level, $l+1$, for the same decomposition procedure. This process is repeated until a predefined recursion depth $\lambda$ is reached, and the final remaining skeleton matrix is set as $D_{\lambda+1}$.
@@
To address this limitation, based on the HBS algorithm [hoFastDirectSolver2012], a more robust strategy uses the geometric coordinates of the underlying data points, which are typically available for matrices defined by a kernel function [coulierInverseFastMultipole2017]. Instead of partitioning indices based on their ordering, the hierarchy is constructed directly from these spatial coordinates, for instance, by employing a quadtree in 2D or an octree in 3D. For matrices that are not strictly generated by a kernel function, precise coordinates are not strictly required for this partitioning; approximate coordinates are sufficient. For example, if the matrix represents a kernel-based system subject to a small perturbation, the coordinates associated with the original kernel function can still be effectively used. This hierarchical structure ensures that matrix blocks corresponding to well-separated spatial clusters are low rank and compressible, whereas blocks corresponding to adjacent points are generally treated as dense diagonal blocks at a given level. The procedure is summarized in Algorithm 3 and illustrated in Figure 1.
@@
nonzero elements.
Proof..
The statement follows from three components in hierarchical matrix compression:
  1. (1)
    Proxy approximation to localize the ID computations (Section 3.1).
  2. (2)
    A level-by-level count of nonzeros and ID costs under uniformly bounded ranks (Section 3.2).
  3. (3)
    Entrywise control of the interpolation matrices produced by strong RRQR (Section 3.3).
Theorem 3.1 states that both the classical preparation cost and the number of nonzero elements in the resulting matrices scale linearly with the matrix dimension $N$. Additionally, the elements of these matrices are uniformly bounded. These properties are essential for constructing sparse matrix oracles and analyzing their quantum costs later in this paper.
In the following subsections, we detail the hierarchical compression procedure that leads to these bounds. We begin by introducing the proxy approximation, which reduces the global compression step to a local one (Section 3.1). Next, we calculate the nonzero count and runtime (Section 3.2), and finally, we explain how the strong RRQR algorithm ensures bounded matrix entries (Section 3.3).
3.1 Proxy Approximation
A major computational cost in the algorithm comes from the compression step. For each block, finding the basis vectors requires information from all other blocks in the same row or column, which is a global operation. If the matrix structure is arbitrary, this is unavoidable. However, for many matrices that come from physical problems, this global work can be replaced with a much cheaper local computation. The main idea, known as proxy approximation [xingInterpolativeDecompositionProxy2020], is to replace the influence of all distant parts of the system with an equivalent set of sources on a small, artificial surface.
As shown in Figure 3, to compress a block corresponding to a cluster of points $X_{0}$, we no longer need to consider all other clusters $Y_{0}$. Instead, we only need to account for its immediate neighbors and a nearby “proxy surface” $\Gamma_{0}$ that separates the local region from the rest of the domain. The field generated by all distant sources can be replicated by an equivalent distribution on this proxy surface. As a result, the compression becomes a local step. The basis for a block is found by interacting it only with its neighbors and the proxy surface, not the entire system. Since the number of points needed for the proxy surface is much smaller and does not depend on the total problem size, this approach can greatly reduce the computational cost.
Figure 3: An illustration of proxy approximation.
3.2 Computational Complexity
In this section, we analyze the storage and computational complexity of constructing the hierarchical representation. The key metrics are determined by the properties of the factor matrices $\{D_{l},L_{l},R_{l}\}$. The total number of nonzero elements in the extended sparse matrix $A_{\text{sp}}$ is of the same order as the sum of nonzeros in these factors. Therefore, an analysis of the storage for these factors provides an effective bound on the overall storage cost of $A_{\text{sp}}$ and allows us to verify that its row and column sparsity is bounded by a constant.
We estimate the total number of nonzero elements, $s$, by summing the contributions from all blocks. The total number of nonzeros in the diagonal blocks is $3^{d}n_{1}^{2}p_{1}$ for $l=1$ and $(6^{d}-3^{d})n_{l}^{2}p_{l}$ for $l>1$. The interpolation matrices and other sparse components also contribute to $s$. Summing these quantities over all levels gives the total number of nonzero elements:
$s$ $=3^{d}n_{1}^{2}p_{1}+\sum\limits_{l=1}^{\lambda}(2k_{l}p_{l}+2k_{l}(n_{l}-k_{l}+1)p_{l}+(6^{d}-3^{d})n_{l}^{2}p_{l})$ (5)
$=3^{d}n_{1}N+(2^{d+1}r^{2}-2r^{2}+4r)(2^{d(\lambda+1)}-2^{d})/(2^{d}-1)+4^{d}(6^{d}-3^{d})r^{2}(2^{d(\lambda+1)}-2^{d})/(2^{d}-1)$
$\sim O(N).$
This linear scaling shows that the extended sparsification reduces the storage requirement from ${O}(N^{2})$ for the original dense matrix to O(N){O}(N).
The computational cost is dominated by the construction of the hierarchical representation, specifically the ID performed at each level. The cost of a rank-$k$ ID on an $m\times n$ matrix is typically ${O}(mn\log k+k^{2}n)$. Summing this over all blocks and levels gives the total complexity for the compression step:
$T\sim\sum\limits_{l=1}^{\lambda}2p_{l}(m_{l}n_{l}\log k_{l}+k_{l}^{2}n_{l}).$ (6)
The key to achieving linear complexity is the use of proxy points, as in Section 3.1. With this approach, the ID for a given block is not performed against all other blocks in the system. Instead, it is only performed against a much smaller matrix formed by its near-neighbors and the proxy points. This reduces the dimension $m_{l}$ in the cost formula from O(N){O}(N) to a small constant, $m$, that depends only on the local geometry. Since the block sizes $n_{l}$ and ranks $k_{l}$ are also bounded by constants ($2^{d}r$ and $r$), the cost to compress a single block becomes constant. The total cost is then simplified to:
$T\sim 2(2^{d}mr\log r+2^{d}r^{3})\sum\limits_{l=1}^{\lambda}p_{l}=2^{d+1}(mr\log r+r^{3})(2^{d(\lambda+1)}-2^{d})/(2^{d}-1)\sim O(N).$ (7)
Therefore, the total computational cost for constructing the hierarchical representation scales linearly with the problem size, $N$.
3.3 Element Bound
For the subsequent analysis, it is important that the entries of the matrices in the HBS representation are bounded. This bound can be guaranteed by constructing the ID using a strong rank-revealing QR (RRQR) algorithm [guEfficientAlgorithmsComputing1996]. We first formalize the definition of this factorization.
Definition 3.2 (Strong RRQR).
Given a matrix $M\in\mathbb{R}^{m\times n}$, we consider partial QR factorizations of the form
$M\Pi=QR\equiv Q\begin{pmatrix}A_{k}&B_{k}\\ &C_{k}\end{pmatrix}$,
where $Q\in\mathbb{R}^{m\times m}$ is orthogonal, $A_{k}\in\mathbb{R}^{k\times k}$ is upper triangular with nonnegative diagonal elements, $B_{k}\in\mathbb{R}^{k\times(n-k)},C_{k}\in\mathbb{R}^{(m-k)\times(n-k)}$, and $\Pi\in\mathbb{R}^{n\times n}$ is a permutation matrix. We call the factorization a strong RRQR if it satisfies
$\sigma_{i}(A_{k})\geq\frac{\sigma_{i}(M)}{q_{1}(k,n)}\quad{\rm{and}}\quad\sigma_{j}(C_{k})\leq\sigma_{k+j}(M)q_{1}(k,n)$
and
$|(A_{k}^{-1}B_{k})_{i,j}|\leq q_{2}(k,n),$
for $1\leq i\leq k$ and $1\leq j\leq n-k$, where $q_{1}(k,n)$ and $q_{2}(k,n)$ are functions bounded by low-degree polynomials in $k$ and $n$.
Gu and Eisenstat [guEfficientAlgorithmsComputing1996] proved that such a factorization can be computed.
Theorem 3.2.
For any given parameter $f\geq 1$, there exists an algorithm that computes a strong RRQR factorization satisfying the conditions in Definition 3.2 with
$q_{1}(k,n)=\sqrt{1+f^{2}k(n-k)}\quad{\rm{and}}\quad q_{2}(k,n)=f.$
Theorem 3.2 ensures that the entries of the interpolation matrix, which corresponds to $A_{k}^{-1}B_{k}$ in the factorization, are bounded by a predefined constant $f$. The entries of the diagonal blocks $D_{l}$ are taken directly from the original matrix $A$ and are therefore also assumed to be bounded. In conclusion, the elements in $A_{\text{sp}}$ are bounded by a constant $c_{\rm sp}$.
4 A Sparsification Based Quantum Approach